# Simul-LLM: A Framework for Exploring High-Quality Simultaneous Translation with Large Language Models

### v1

@@ 7.2 Exploration of SimulMT LLMs with Proposed Prompt Structure
In Table 1, we also provide a breakdown of the performance of our exploration of SimulMT LLMs with our proposed prompt structure in Section 5.2 that carefully manages source context availability for target translation generation. Two models are employed for this exploration, one fine-tuned for wait-3 inference, and another fine-tuned for wait-7 inference that produces noticeable better quality translations than the first.
While SimulMT LLMs are a more promising approach in achieving higher translation quality than NMT LLMs due to more direct task-specific fine-tuning and better context alignment, our experiment results suggest that, for the time-being, the performance of SimulMT LLMs is not advantageous to NMT LLMs adapted to SimulMT (and varies from slightly worse to barely better translation quality compared with the classical non-LLM SimulMT baseline). This is not completely unexpected as NMT LLMs have been optimized heavily in recent years whereas the exploration of SimulMT LLMs has just started. We provide some analysis below that point out several possible reasons for this observed performance gap and call for additional community efforts to investigate further.
First, due to computational constraints we were unable to fine-tune for an entire epoch of the training dataset (only 2M random samples out of 5M), which represents a major loss of lessons and a lack of rigorous wait-k curriculum for the fine-tuned model. Second, it is possible that the fine-tuning hyperparameters are ill-suited for this particular prompt. We consider this likely to be the most influential issue on our observed results, given the drastic differences between the original and expanded datasets. Third, at least one other work related to NMT LLMs (Chen et al., 2023) has demonstrated that relative positional embeddings can cause issues via attention dilution that ends up being unhelpful, suggesting that distancing the source context, running target hypothesis, and the current translation step hypothesis can be unexpectedly problematic. We posit that our proposed Simul-LLM can be leveraged to verify the above reasons.
@@ 7.3 Higher Wait-k Generalizability Comparisons
It is well documented that in typical SimulMT systems, training or fine-tuning with a slightly higher wait-k than is intended during inference can boost translation quality and generalizability across slightly lower wait-k (Ma et al., 2019). While this likely applies to SimulMT LLMs, no existing work has validated that this behavior persists. We provide a brief comparison of two SimulMT LLMs fine-tuned via wait-3 and wait-7 context levels in Table 2. The results in the table demonstrate that, generally, the expected behavior does hold, with all LLMs fine-tuned in wait-7 outperforming their corresponding wait-3 models for the same inference wait-k. We leave validating additional, previously understood SimulMT principles in SimulMT LLMs to future work.
@@ 8 Conclusion
library as well as with SimulEval, the preeminent SimulMT evaluation framework. In addition to introducing Simul-LLM, we employ this framework to explore adapting existing NMT LLMs to SimulMT and propose an expanded fine-tuning dataset and alternative prompt structure to those employed in typical NMT fine-tuning that we believe better replicates inference-time behavior. Moreover, we validate some classically understood SimulMT concepts concerning wait-k scheduling and examine the behavior of SimulMT LLMs during fine-tuning. Our proposed Simul-LLM framework enables multiple lines of future work that can be carried out to understand and optimize LLMs for SimulMT, and it will likely be a useful tool for the research community.

### v4

@@ 7.2 Exploration of SimulMT LLMs with Proposed Prompt Structure
In Table 1, we also provide a breakdown of the performance of our exploration of SimulMT LLMs with our proposed prompt structure in Section 5.2 that carefully manages source context availability for target translation generation. Two models are employed for this exploration, one fine-tuned for wait-3 inference, and another fine-tuned for wait-7 inference that produces noticeably better quality translations than the first. In addition, we provide results for the fine-tuned wait-7 models after fine-tuning on 2M samples and 5M samples of the fine-tuning dataset, with the fine-tuned wait-3 model having fine-tuned for 2M samples. This separation is due to computational constraints on our side, as we wanted to provide further comparisons in Section 7.3 and Table 2 at various evaluation wait-k values but lacked the compute to finish fine-tuning the wait-3 model.
While SimulMT LLMs are a more promising approach in achieving higher translation quality than NMT LLMs due to more direct task-specific fine-tuning and better context alignment, our experimental results suggest that, for the time-being, the performance of SimulMT LLMs is not guaranteed to be advantageous compared to NMT LLMs adapted to SimulMT. As seen in Table 1, fully fine-tuned SimulMT LLMs outperformed their NMT LLM counterparts on en-es but failed to do so on en-de. This is not completely unexpected as NMT LLMs have been optimized heavily in recent years whereas the exploration of SimulMT LLMs has just started. We provide some analysis below that points out a few possible reasons for this observed performance gap and we call for additional community efforts to investigate further.
First, it is possible that the fine-tuning hyperparameters are ill-suited for this particular prompt. We consider this likely to be the most influential issue on our observed results, given the drastic differences between the original and expanded datasets (the fine-tuning hyperparameters were united for both fine-tuning tasks). Second, at least one other work related to NMT LLMs (Chen et al., 2023) has demonstrated that relative positional embeddings can cause issues via attention dilution that ends up being unhelpful, suggesting that distancing the source context, running target hypothesis, and the current translation step hypothesis can be unexpectedly problematic. We posit that our proposed Simul-LLM can be leveraged to verify the above reasons.
@@ 7.3 Higher Wait-k Generalizability Comparisons
It is well documented that in typical SimulMT systems, training or fine-tuning with a slightly higher wait-k than intended during inference can boost translation quality and generalizability across slightly lower wait-k (Ma et al., 2019). While this likely applies to SimulMT LLMs, no existing work has validated that this behavior persists. We provide a brief comparison of two SimulMT LLMs fine-tuned via wait-3 and wait-7 context levels in Table 2. The results demonstrate that, generally, the expected behavior does hold, with all LLMs fine-tuned in wait-7 outperforming their corresponding wait-3 models for the same inference wait-k with up to a 2.6 BLEU improvement. We leave validating additional, previously understood SimulMT principles in SimulMT LLMs to future work.
7.4 Discussion on Other Works
We are aware of two other concurrent efforts at applying LLMs to simultaneous translation (Koshkin et al., 2024; Wang et al., 2023). The former focuses on high-resource settings (i.e. order-of-magnitude larger language models than we aim for in this work) and additionally touches on comparisons with GPT-4 and speech-to-text simultaneous translation efforts. Contrastingly, the latter work aligns with ours, approaching the problem of LLM simultaneous translation in a low-resource setting. Their work, however, is limited to a focused contribution: a variation of longest-common prefix (LCP) that they call relaxed agreement LCP (RALCP). Neither host open-source software to test their changes and both only test Llama-based models in addition to arguably shallower explorations of this subject matter. Nonetheless, we intend to add support for end-to-end speech-to-text translation soon and have already added support for RALCP, which we provide brief results for in our Appendix. Both works validate our choice to engage with MuST-C as a dataset, as both works also employ it.
@@ 8 Conclusion
library as well as with SimulEval, the preeminent SimulMT evaluation framework. In addition to introducing Simul-LLM, we employ this framework to explore a wide range of topics in the LLM SimulMT space. Our proposed Simul-LLM framework enables multiple lines of future work that can be carried out to understand and optimize LLMs for simultaneous translation, and it will likely be a useful tool for the research community.
9 Limitations
This work is focused on providing an open-source framework for an interesting application of LLMs. While we propose a few novel techniques for engaging with LLMs in this application space that we support in Simul-LLM, the work is largely exploratory and we cannot assert that our proposed practices are guaranteed to work best. Additionally, the framework itself does not currently support speech-to-text exploration (although augmentation to this end should not be difficult), a popular simultaneous translation track. All of these points serve to slightly limit the impact of our contribution.
10 Acknowledgements
This research was supported, in part, by the National Science Foundation grants 2223483 and 2223484.
Appendix A Appendix
A.1 Licensing Information
Fairseq (Ott et al., 2019) is MIT-licensed. SimulEval is licensed via CC BY-SA 4.0. MuST-C, one of the premiere speech-to-text datasets across many language pairs, is licensed under CC BY-NC-ND 4.0. Simul-LLM itself is MIT-licensed.
A.2 Training, Fine-Tuning, and Evaluation Hyperparameters and Hardware Details
All classical models were trained on two NVIDIA 32GB V100s and validated on a single V100. All LLMs were fine-tuned via PEFT Mangrulkar et al. (2022) on a single NVIDIA 40GB A40 in bfloat16 and evaluated on a single V100 in float32. Simul-LLM seamlessly integrates with SimulEval (Ma et al., 2020) for the purpose of these evaluations. Classical transformer baselines were trained via Fairseq (Ott et al., 2019), an easily extensible sequence to sequence modeling toolkit written in PyTorch.
Classical models were trained with typical hyperparameters provided in Fairseq examples. All LLMs were fine-tuned with identical hyperparameters, employing a constant learning rate of 3e-4 and were optimized via Adam with around 4K warmup updates and batch sizes of 40 samples. LoRA adapter parameters were an of 16 and a value of 64, resulting in a total of around 40M added parameters during fine-tuning, with a dropout value of 0.1. For fair comparison, classical models were of a similar, although not quite identical, size in terms of parameter count (around 46M parameters).
Additionally, all LLMs were fine-tuned while quantized with NormalFloat4 (nf4) quantization. A small performance boost was observed when removing this quantization during inference, so all models did not engage with nf4 quantization during inference.
A.3 Estimated Computational Costs
We estimate that, normalizing for 1 GPU, building Simul-LLM cost roughly 26 GPU days of experimentation and approximately 24 GPU days to generate the results in this paper.
Model | CA LAAL (ms) |
Monotonic Transformer Wait-5 | 1325 |
Wait-7 Fine-tuning LLM | 3237 |
A.4 Computational Latency of LLMs for SimulMT
While we only cover results for lagging behavior (via LAAL) in the main body of this paper, it is worth acknowledging that employing an LLM for SimulMT incurs additional computational latency costs compared to a more classical solution. These costs can be characterized in a number of ways, but we choose to employ a latency-based version of LAAL inspired by simultaneous speech translation (SimulST), where we replace word-based lagging measurements with estimated acoustic latency in milliseconds. This allows us to simply add the computational cost of each generation step to the inherent, estimated acoustic latency (we employ 360 ms of inherent latency, equivalent to around 166 words/min). Adding this inherent latency is akin to treating the SimulMT system as though an oracle, zero-latency transcription agent has been prepended to it and is critical to ensuring that LAAL operates correctly, as a metric. Brief latency results can be observed in Table 3. For further details, we refer readers to the original LAAL publication (Papi et al., 2022).
Decoding Scheme | en-de | en-es |
Chunk SBS | 25.8 (5.6) | 27.0 (5.9) |
Chunk SBS w/ RALCP | 26.1 (5.3) | 28.1 (5.5) |
A.5 RALCP and Policy Differences
While Simul-LLM supports RALCP (Wang et al., 2023), our application differs slightly from the specification of the original implementation. Unlike the authors of Wang et al. 2023, we do not engage in consecutive read decisions unless it is to maintain a k-lagging factor as specified by an inference wait-k value. This ensures, after a given set of outputted chunks via RALCP implemented on top of SBS (Zheng et al., 2020), that expected context limitations are kept. This contrasts with the original implementation of RALCP on top of a more relaxed read-write schedule that does not always attempt to maintain a k-lagging factor after that lagging factor is initially established. Additionally, we assume that the authors applied a word-based simultaneous approach, as this is not specified in their original work.
Nonetheless, RALCP, which is fundamentally a form of beam rescoring, does improve translation quality based on our empirical results. We include a brief comparison in Table 4, where we observe that the addition of RALCP seems to result in small to moderate improvements in BLEU score with roughly equivalent lagging behavior (slight reductions in our tests).
A.6 Supporting Various LLMs
Simul-LLM’s easy to extend fine-tuning and evaluation framework supports Falcon, Llama, and Mistral-based models. While our explorations in this paper are rooted in Falcon-based models, we provide a brief set of results for other LLMs in Table 5. Generally, translation quality appeared similar across models, although the Llama-based model struggled. BLEU degradation for Llama models seemed mostly linked to degenerate output where the model would get stuck in low-quality generation (e.g. continuously outputting the same token).
Foundational Model | BLEU (LAAL) |
Falcon 7B LLM | 30.06 (3.95) |
Mistral 7B LLM | 30.25 (3.66) |
Llama-2 7B LLM | 20.18 (3.84)444Some issues were observed during fine-tuning the Llama-based translation agent that seems to have resulted in poorer than expected translation quality. We are actively working to improve these results within the Simul-LLM framework. |